OpenStudy (anonymous):

8-10 deal with the ambiguous SSA case. for each find all possible solutions and sketch the triangle in each case. 8. A=50 degrees, a=15, b=12 9. A=50 degreees, a=10, b=15 10. A=47 degrees, a=1.5 b=2

OpenStudy (cwtan):

15/sin50=12/sin b sinb=12sin50/15 =0.6128.... b=37.7948.......(37'47'41.34) or b= 180- 37.7948.... =142.2052......(142'12'18.6) Using the same way to solve other problem